OpenStudy (anonymous):

need help solving for a.. A=1/2 h(a/b) how do you find the solution?need steps

OpenStudy (dumbcow):

so you need to get a by itself on right side \[A = \frac{h*a}{2b}\] \[\rightarrow 2b*A = h*a\] \[\rightarrow \frac{2b*A}{h} = a\]